A partnership is considering possible liquidation because one of the partners (Bell) is personally insolvent. Profits and losses are divided on a 4:3:2:1 basis, respectively. Capital balances at the current time are

Bell, capital $ 83,500
Hardy, capital 69,000
Dennard, capital 16,000
Suddath, capital 93,000

Bell’s creditors have filed a $34,000 claim against the partnership’s assets. The partnership currently holds assets of $430,000 and liabilities of $168,500. If the assets can be sold for $255,000, what is the minimum amount that Bell’s creditors would receive?

Carney, Pierce, Menton, and Hoehn are partners who share profits and losses on a 4:3:2:1 basis, respectively. They are beginning to liquidate the business. At the start of this process, capital balances are

Carney, capital $ 78,000
Pierce, capital 32,400
Menton, capital 61,000
Hoehn, capital 25,400

Which of the following statements is true?

Multiple Choice

  • Carney will be the last partner to receive any available cash.

  • The first available $7,400 will go to Hoehn.

  • The first available $10,200 will go to Menton.

  • Carney will collect a portion of any available cash before Hoehn receives money.
Answer #1

1. Loss on sale of assets = $430000 - $255000

= $175000

Share of Bill = $175000*4/10

= $70000

Opening balance of Bill = $83500

After adjusting loss = $83500 - $70000

= $13500

Hence his creditors will receive minimum $13500.

2. Who will receive cash first will be decided on the basis of capital amount divided by ratio.

Carney = $78000/4 = $19500

Pierce = $32400/3 = $10800

Menton = $61000/2 = $30500

Hoehn = $25400/1= $25400

We have to make all capital per ratio equal , hence since Menton has the highest capital per ratio , he will first receive cash unless it becomes equivalent to hoehn.

= $30500 - $25400

= $5100 per ratio

Now we have to multiply it by ratio of monten

= $5100*2 = $10200

Hence after receiving $10200 , montens capital per ratio will become equal to hoehn.

So the answer is option C.